# Enhancing Brain Source Reconstruction through Physics-Informed 3D Neural Networks

### v1

@@ VI-A Limitations & Future Work
The quality of the data prior highly depends on the simulation used to train the model. While we used a diverse simulation setting, varying size and number of the sources, we impose a Gaussian prior on the source shape, which could be replaced by more realistic assumptions to refine the prior in future work. Additionally, incorporating known correlation structures between sources could further improve the learned prior.
Furthermore, our focus on the spatial aspects of source localization neglects temporal dynamics. As the information contained in a single measurement is limited, incorporating the time domain could further improve the accuracy, especially in distinguishing multiple active sources. However, the proposed hybrid approach is general and the core idea of combining pseudo-inverse with deep learning can be adapted to also include the time domain, which is a direction we plan to pursue in future work.
Like many advanced learning approaches, the proposed 3D-PIUNet presents challenges in interpretability. Neural networks, while highly effective, can sometimes be seen as “black boxes”. In future research, we aim to integrate the recent advances on neural network explainability [70] to enhance the transparency and reliability of neural network-based source localization methods.

### v2

@@ VI-A Limitations & Future Work
The effectiveness of our method depends on the quality of the data prior, which is shaped by the simulations used during training. While we employed a diverse set of simulations, varying the size and number of sources, our approach assumes a Gaussian prior on the source shape. This assumption, while effective for controlled evaluations, may not fully capture the diversity and complexity of real neural activations, potentially limiting performance in highly heterogeneous clinical datasets. Future work could explore more biologically informed priors, such as those based on functional parcellations [17], to enhance realism. Additionally, incorporating known correlation structures between sources, as common in neural mass models [76, 6], could further refine the learned prior and improve source reconstruction in real-world settings.
Furthermore, our focus on the spatial aspects of source localization neglects temporal dynamics. Given the inherently ill-posed nature of EEG/MEG source localization, using temporal information could enhance accuracy, particularly in resolving multiple simultaneously active sources. However, the proposed hybrid approach is general and the core idea of combining pseudo-inverse with deep learning can be adapted to also include the time domain, which is a direction we plan to pursue in future work. This extension is particularly relevant for clinical applications that focus on the temporal dynamics of neural events, such as brain-computer interfaces.
Like many advanced learning approaches, the proposed 3D-PIUNet presents challenges in interpretability. Neural networks, while highly effective, can sometimes be seen as “black boxes”. This can pose challenges in clinical adoption, where transparency and trust in model decisions are essential. In future research, we aim to integrate the recent advances on neural network explainability [62] to improve the transparency and reliability of neural network-based source localization methods.